How many molecules of hydrochloric acid (HCI) are required to react with 10.0

grams of zinc (Zn)?


1
Expert's answer
2022-04-19T03:11:58-0400

Zn + 2HCl = ZnCl2 + H2

n(Zn) = m(Zn) / Mr(Zn) = 10.0 g / 65.38 g/mol = 0.153 mol

1 mol Zn - 2 mol HCl

0.153 mol Zn - x mol HCl

x = 0.153 × 2 = 0.306 mol

N(HCl) = n(HCl) × NA = 0.306 × 6.02×1023 = 1.84×1023